# Towards Autonomous Indoor Parking: A Globally Consistent Semantic SLAM System and A Semantic Localization Subsystem

### v1

@@ III-C Map-Based Localization Subsystem
$\mathcal{G}_{\text{SF-Loc}}=\{\{N_{pose}^{i}\},\{E_{odo}^{i,j}\},\{E_{ICP}^{i}\}\}.$ (10)
There is only one type of pose node $\{N_{pose}^{i}\}$, which is the same as the Pose Node introduced in Section III-A1. The Odometry Error Term (OET) is described in Section III-A3. Additionally, some nodes are constrained by Semantic ICP Error Terms, which are based on the semantic ICP registration.
@@ IV EXPERIMENTS
(a) Dataset 1
(b) Dataset 2
Figure 7: The global maps built by GCSLAM on Dataset 1 and Dataset 2.
@@ IV-A1 Tracking
As shown in Table I, GCSLAM demonstrates significantly lower absolute trajectory error than VIW [18] and Ground-Fusion [16]. ORB-SLAM2 [11] and DM-VIO [15] fail due to the absence of wheel encoders. In indoor parking lots, cameras can not offer reliable data because of the complex lighting conditions, while wheel encoders can provide more accurate data, as the ground is flat and tire slippage is negligible.
@@
Dataset SWE(cm) AE(cm)
Dataset 1 0.044 2.146
Dataset 2 0.492 0.776
@@ IV-A2 Mapping
We evaluate mapping performance by two proposed metrics: Slot Width Error (SWE) and Adjacent Error (AE). SWE denotes the difference between the average parking slot width in the global map and the actual slot width, representing the gap between the global map and the real world. AE equals to the average $||E_{adj}||$ calculated for each pair of adjacent slots. $||E_{adj}||$ is defined in Equation 5. AE represents the distance between adjacent slots, which is supposed to be zero. Fig. 7 shows the global maps. Table II presents the accuracy results of GCSLAM mapping, demonstrating the precision of our algorithm.
@@ IV-B Evaluation on SF-Loc
As SF-Loc is based on known map localization, we focus on evaluating its accuracy within the global map. Therefore, We use the GCSLAM trajectory as the ground truth. The global map we use is shown in Fig. 7. The Semantic ICP is our registration algorithm introduced in Section III-C. As shown in Table III, semantic ICP significantly improved accuracy compared to vanilla ICP. Moreover, the SF-Loc outperformed semantic ICP results attributed to the semantic fusion factor graph.
@@ V CONCLUSION
In this paper, we present GCSLAM, a novel system for indoor parking tracking and mapping. GCSLAM incorporates an innovative factor graph and novel error terms, enabling robust and high-precision mapping in complex parking environments. Additionally, we develop a map-based localization subsystem SF-Loc. SF-Loc fuses registration results and odometry poses based on a novel factor graph, effectively enhancing localization precision. We validate our algorithm through real-world datasets, demonstrating the effectiveness and robustness of our system.

### v2

@@ III-C Map-Based Localization Subsystem
$\mathcal{G}_{\text{SF-Loc}}=\{\{T_{node}^{i}\},\{E_{odo}^{i,j}\},\{E_{ICP}^{i}\}\}.$ (10)
There is only one type of pose node $\{T_{node}^{i}\}$, which is the same as the Pose Node introduced in Section III-A1. The Odometry Error Term (OET) is described in Section III-A3. Additionally, some nodes are constrained by Semantic ICP Error Terms, which are based on the semantic ICP registration.
@@ IV EXPERIMENTS
Figure 7: The global maps constructed by GCSLAM and MOFISSLAM [3] on Dataset 1 and Dataset 2. $\text{MOFISSLAM}^{*}$ [3] is a reproduction of our own implementation. For fairness, we incorporated wheel-related constraints into the process.
@@ IV-A1 Tracking
As shown in Table I, GCSLAM demonstrates significantly lower absolute trajectory error than ORB-SLAM3 [23], VIW [19] and Ground-Fusion [17]. This is because our semantic-constrained factor graph effectively provides more stable and global semantic constraints, resulting in minimal cumulative drift, better accuracy and global consistency. DM-VIO [16] fail due to the absence of wheel encoders. In indoor parking lots, cameras can not offer reliable data because of the complex lighting conditions, while wheel encoders can provide more accurate data, as the ground is flat and tire slippage is negligible.
@@
Methods Dataset 1 Dataset 2
SWE(cm) $\downarrow$ AE(cm) $\downarrow$ SWE(cm) $\downarrow$ AE(cm) $\downarrow$
$\text{MOFISSLAM}^{*}$ [3] 0.703 4.224 1.729 4.279
Ours (GCSLAM) 0.044 2.146 0.492 0.776
@@ IV-A2 Mapping
We evaluate mapping performance by two proposed metrics: Slot Width Error (SWE) and Adjacent Error (AE). SWE denotes the difference between the average parking slot width in the global map and the actual slot width, representing the gap between the global map and the real world. As shown in the Fig. 7, we compared GCSLAM and $\text{MOFISSLAM}^{*}$ [3] on our datasets. Table II presents the accuracy results of both methods, demonstrating the precision of our algorithm. GVET improves the global consistency of GCSLAM, while the unstable slots filter and slot updater in global slot management module allow GCSLAM to better adapt to noise. Due to the absence of GVET and Global Slot Management, $\text{MOFISSLAM}^{*}$ [3] exhibits misalignment and distortion in large-scale, complex parking lots.
@@ IV-B Evaluation on SF-Loc
As SF-Loc is based on known map localization, we focus on evaluating its accuracy within the global map. Therefore, We use the GCSLAM trajectory as the ground truth. The global map we use is shown in Fig. 7. The Semantic ICP is our registration algorithm introduced in Section III-C. As shown in Table III, vanilla ICP fails due to convergence failure in the sparse map area. Compared to vanilla ICP, semantic ICP significantly improved accuracy. Moreover, the SF-Loc outperformed semantic ICP results attributed to the semantic fusion factor graph.
@@ V CONCLUSION
In this paper, we present GCSLAM, a novel system for indoor parking tracking and mapping. GCSLAM incorporates a semantic-constrained factor graph and novel error terms, enabling robust and high-precision mapping in complex parking environments. Additionally, we develop a map-based localization subsystem SF-Loc. SF-Loc fuses registration results and odometry poses based on a novel factor graph, effectively enhancing localization precision. We validate our algorithm through real-world datasets, demonstrating the effectiveness and robustness of our system.